# Localized statistics decoding: A parallel decoding algorithm for quantum low-density parity-check codes

### v1

@@ 3 $\mathbf{\lambda}$: fault node soft information vector
4 $\mathbb{I}\coloneqq[\,]$: list of invalid clusters
5 $\mathbb{V}\coloneqq[\,]$: list of valid clusters
@@
7       $\rm cl_{i}$ = create_cluster($s_{i}$)
8       $\mathbb{I}$.add($\rm cl_{i}$)
@@ 7       $\rm cl_{i}$ = create_cluster($s_{i}$)
14      for $(\rm cl_{i},\rm cl_{\ell})\in(\mathbb{I}\cup\mathbb{V})$ do
15             merged = check_collision$(\rm cl_{i},\rm cl_{\ell})$
@@
17                   $\rm cl_{i\cup\ell}=$ merge_clusters$(\rm cl_{i},\rm cl_{\ell})$
18                   $\mathbb{I}$.remove($\{\rm cl_{i},\rm cl_{\ell}\}$)
19                   $\mathbb{V}$.remove($\{\rm cl_{i},\rm cl_{\ell}\}$)
20                   $\mathbb{I}$.add($\rm cl_{i\cup\ell}$)
@@ 17                   $\rm cl_{i\cup\ell}=$ merge_clusters$(\rm cl_{i},\rm cl_{\ell})$
24             valid = cl.check_validity($\mathbf{s}_{[\rm cl]}$)
@@
30 for cl $\in\mathbb{V}$  do
[…]
@@ 4.2 On-the-fly elimination
To grow and merge clusters, multiple smaller steps are necessary. As detailed above, these steps include identifying fault nodes/column indices of the decoding matrix $H$ by which the invalid clusters will grow and determining whether an added fault node will lead to two or more clusters merging into a single one – see Definition 4.4. For simplicity, we first describe the case of sequential cluster growth. Our on-the-fly procedure can analogously be applied in a parallel implementation, cf. Appendix B.
Let $C_{i}$ be an active cluster, that is, $\left(H_{[C_{i}]},\mathbf{s}_{[C_{i}]}\right)$ does not define a solvable decoding problem as $s_{[C_{i}]}\notin\textsc{image}(H_{[C_{i}]})$. To grow cluster $C_{i}$, we consider candidate fault nodes $v_{j}\in\Lambda(C_{i})$ – fault nodes not already in $C_{i}$ but connected to check nodes on its boundary $\beta(C)$, see Eq. 4. The candidate fault node with the highest probability of being in error according to the soft information vector $\mathbf{\lambda}\in\mathbb{R}^{n}$ is selected. Once $v_{j}$ has been chosen, we check whether its neighboring detector nodes are boundary nodes of any other (valid or invalid) clusters i.e., we check for collisions, cf Definition 4.4. If this is not the case, we proceed as follows. We now assume that the active cluster $C_{i}$ described by sub-matrix $H_{[C_{i}]}$ that has a PLU factorization of the form
@@
where $P_{i},L_{i},U_{i}$ are as in Eq. 6. Adding a fault node $v_{j}$ to the cluster is equivalent to adding a (sparse) column vector $\mathbf{b}$ to $H_{[C_{i}]}$, i.e.,
@@
$\begin{pmatrix}\begin{array}[]{c|c}U_{i}&L_{i}^{-1}P_{i}\mathbf{b}_{i}\\ 0&\mathbf{b}_{i*}\end{array}\end{pmatrix},$ (9)
where $\mathbf{b}_{i}$ is the projection of $\mathbf{b}$ onto the detectors/row coordinates that are enclosed by $C_{i}$. Similarly, $\mathbf{b}_{i*}$ is the projection onto detector coordinates not enclosed by $C_{i}$. Importantly, applying the operators $L_{i}^{-1}$ and $P_{i}$ does not affect the support of $\mathbf{b}_{i}$ and $\mathbf{b}_{i*}$, as both these operators act solely on the support of $C_{i}$. Combining this with Eq. 9, we note that to complete the PLU factorization of $H_{[C_{i}\cup\{v_{j}\}]}$ only $\mathbf{b}_{i*}$ has to be reduced, which has a computational cost proportional to its weight – crucially only a small constant for bounded LDPC matrices $H$.
@@
Suppose that the selected fault node $v_{j}$ by which the cluster $C_{i}$ is grown is connected to a check node in the boundary $\beta(C_{\ell})$, with $C_{i}\neq C_{\ell}$. Let $\mathbf{b}$ be the column of $H$ associated with the fault node $v_{j}$. Re-ordering its coordinates if necessary, we we can write $\mathbf{b}$ as $(\mathbf{b}_{i},\mathbf{b_{\ell}},\mathbf{b_{*}})$ where $\mathbf{b}_{i},\mathbf{b_{\ell}}$, and $\mathbf{b_{*}}$ are the projections of $\mathbf{b}$ on the row coordinates contained in $C_{i}$, $C_{\ell}$, and neither of them, respectively. Thus, using a block matrix notation, for the combined cluster $C_{i}\cup C_{\ell}\cup\{v_{j}\}$, we have
@@
$\begin{pmatrix}\begin{array}[]{c|c|c}U_{i}&0&L_{i}^{-1}P_{i}\mathbf{b}_{i}\\ \hline\cr 0&U_{\ell}&L_{\ell}^{-1}P_{\ell}\mathbf{b}_{\ell}\\ \hline\cr 0&0&\mathbf{b}_{*}\end{array}\end{pmatrix}.$ (11)
@@
Here, by slight misuse of notation, we group the non-zero rows of $U_{m}$ in the index set $(m,\bullet)$, and its zero rows in the set $(m,\bot)$; we regroup the coordinates of vector $\mathbf{u}$ accordingly. We remark that the row sets $\bullet$ and $\bot$ are distinct from the row set * identified when writing $\mathbf{b}$ as the combination of its projection onto row coordinates enclosed by $C_{i}$ and outside it. By identifying the appropriate row sets for the clusters $C_{i}$, $C_{\ell}$ and the fault node $\{v_{j}\}$ as detailed in Eq. 17, we can construct a block-swap matrix to bring $H_{[C_{i}\cup C_{\ell}\cup\{v_{j}\}]}$ into the form
$\begin{pmatrix}\begin{array}[]{c|c|c}U_{i,0}&0&\mathbf{u}_{i,\bullet}\\ \hline\cr 0&U_{\ell,\bullet}&\mathbf{u}_{\ell,\bullet}\\ \hline\cr 0&0&\mathbf{u}_{i,\bot}\\ \hline\cr 0&0&\mathbf{u}_{\ell,\bot}\\ \hline\cr 0&0&\mathbf{b}_{*}\end{array}\end{pmatrix},$ (18)
and similarly for its PLU factors. Since the matrix in Eq. 18 has the same form as the one in Eq. 8, the algorithm can proceed from this point onward as in the case of the addition of a single fault node to a cluster. In conclusion, via a swap transformation, we can effectively reduce the problem of merging two clusters to the problem of adding a single fault node to one cluster.
@@ 4.3 Numerical decoding simulations
  • Qubits initialized in |$\ket{0}(\ket{+}$) are flipped to |$\ket{1}$ ($\ket{-}$) with probability $p$.
  • The measurement result of an $X$/$Z$ basis measurement is flipped with probability $p$.
@@
The syndrome extraction circuit for the family of HGP codes presented in Section C.2 and results presented in Section 2.3 is obtained from the minimum edge coloration of the Tanner graphs associated to the respective parity check matrix, see Ref. [13] for details. In particular, we generate associated Stim files of $r=12$ noisy syndrome extractions using a publicly available implementation of the aforementioned coloration circuit by Pattison [55]. In this case, the standard circuit-level noise model described at the beginning of this section is employed. We decode $X$ and $Z$ detectors separately using a $(3,1)-$overlapping window decoder. That is, for each decoding round, the decoder obtains the detection events for $w=3$ syndrome extraction cycles and computes a correction for the entire window. However, it only applies the correction for a single $(c=1)$ syndrome extraction cycle, specifically the one that occurred the furthest in the past. For more details on circuit-level overlapping window decoding, see Ref. [56]. We have chosen $w=3$ as this was the value used in Ref. [13]. Note that it is possible that (small) decoding improvements could be observed by considering larger values of $(w,c)$ for the overlapping window decoder.
@@ 4.4 Parallel algorithm
We propose a parallel version of the LSD algorithm (P-LSD) in Algorithm 2 that uses a parallel data structure, inspired by Refs. [58, 59], to minimize synchronization bottlenecks. We discuss the parallel algorithm in more detail in Appendix B. There, we derive a bound on the parallel depth of P-LSD, that is, roughly the maximum overhead per parallel resource of the algorithm. We show that the depth is $O(\operatorname{polylog}(n)+\kappa^{3})$ in the worst-case, where $n$ is the number of vertices of the decoding graph and $\kappa$ is the maximum cluster size. A crucial factor in the runtime overhead of P-LSD is given by the merge and factorization operations. We contain this overhead by (i) using the parallel union-find data structure of Ref. [58] for cluster tracking and (ii) using parallel on-the-fly elimination to factorize the associate matrices. If we assume sufficient parallel resources, the overall runtime of parallel LSD is dominated by the complexity of computing the factorization for the largest cluster. To estimate the expected overhead induced by the cluster sizes concretely, we (i) investigate analytical bounds and (ii) conduct numerical experiments to analyze the statistical distribution of clusters for several code families, see Appendix A.
@@ 7 Acknowledgments
This work was done in part while L.B. was visiting the Simons Institute for the Theory of Computing. L.B. and R.W. acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 101001318) and Millenion, grant agreement No. 101114305). This work was part of the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus, and has been supported by the BMWK on the basis of a decision by the German Bundestag through project QuaST, as well as by the BMK, BMDW, and the State of Upper Austria in the frame of the COMET program (managed by the FFG).
J. R. acknowledges support from EPSRC grants EP/T001062/1 and EP/X026167/1. The Berlin team has also been funded by BMBF (RealistiQ, QSolid), the DFG (CRC 183), the Munich Quantum Valley, the Einstein Research Unit on Quantum Devices, the Quantum Flagship (PasQuans2, Millenion), and the European Research Council (ERC DebuQC). For Millenion and the Munich Quantum Valley, this work is the result of joint-node collaboration.
@@ Appendix A Clusters on decoding graphs
To model an error distribution on the fault graph $\mathcal{F}(H)$, we assume that each fault node is occupied randomly with probability $p$ and empty with probability $1-p$. Hence, for $n$ faults, we expect $np$ faults to be occupied, and $n(1-p)$ to be empty. This setup defines a site percolation problem, which involves studying the distribution of groups of neighboring occupied faults, called clusters. We call any instance of the site percolation problem an error defined on the fault graph $\mathcal{F}(H)$, and we expect clusters to have varying sizes and shapes depending on both the fault graph connectivity and the flip probability $p$.
@@
    the number $\nu$ of clusters of the error;
@@
More broadly, the expected performance depends on the average size $\kappa_{\alpha}$ of the clusters of the error.
If we assume enough parallel resources, LSD could run on $\nu$ cores, in $O(\kappa^{3})$ time in the worst-case and in $O(\kappa_{\alpha}^{3})$ time on average, to solve the local decoding problems, see Appendix B. Following this reasoning, we collect statistical data on these three quantities $\nu,\kappa$ and $\kappa_{\alpha}$, for a given noise model of interest. We study the expectation value of $\nu,\kappa$ and $\kappa_{\alpha}$ for:
@@
In Figure 3, we report the statistics for the fault graph obtained from the circuit-level noise simulation of the $\llbracket 144,12,12\rrbracket$ code [12]. As we can see, for noise values below the threshold, there is no statistical difference between the two different cluster distributions, indicating that the clusters found by BP+LSD are of minimal size and hence (close to) optimal.
@@ A.1 Analytic cluster bounds
Figure 7: Distribution of the cluster size statistics of the $\llbracket 144,12,12\rrbracket$ bivariate bicycle code of Ref. [12] under circuit-level noise with strength $p$. Markers show the mean of the distribution while shapes are violin plots of the distribution obtained from $10^{5}$ samples. Yellow distributions show statistics for the optimal factorization while the blue distributions show statistics for the factorization returned by BP+LSD. The dashed line represents the expected average cluster size for the Bethe lattice, $\kappa_{\alpha}$, for $\theta=139$. Markers and distributions are slightly offset from the actual error rate to increase readability.
@@
If $H$ is an LDPC matrix with constant row and column weight $r$ and $c$, respectively, the associated fault graph $\mathcal{F(H)}$ is a regular graph of degree $\theta=c(r-1)$, with ${\theta\cdot n}/{2}$ edges.
Let $n_{s}^{p}(v)$ be the probability that the node $v$ belongs to a particular cluster of size $s$, an $s$-cluster. In general, this takes the form
$p^{s}(1-p)^{|\mathrm{perimeter}|}$ (19)
where the perimeter of a cluster is the set of non-occupied nodes adjacent to at least one node in the cluster. Summing over the number $a_{s,t}(v)$ of clusters of size $s$ and their perimeter $t$, we find that the probability of belonging to an $s$-cluster can be written as a sum of the form
$n_{s}^{p}(v)=\sum_{t}a_{s,t}(v)p^{s}(1-p)^{t}.$ (20)
On our percolation problem, by assumption, the probability that a given node is occupied is $p$. Moreover, if a node is occupied, it belongs to a cluster of size $\tilde{s}$ for some $\tilde{s}=1,\dots,n$, where $n$ is the number of nodes in the graph. Hence, using Eq. 20, for any node in the graph, it holds that
$\sum_{s}n_{s}^{p}(v)s=p,$ (21)
i.e., the probability $n_{s}^{p}(v)$ that a node belongs to an $s$-cluster times the number of nodes/choices for that cluster.
A quantity of interest for decoding is an estimate of the size of the cluster a random node belongs to. In other words, the probability $w_{s}$ that a node is occupied and belongs to an $s$-cluster. Combining Eq. 20 and Eq. 21,
@@
we find that the average cluster size is given by
@@
For the Bethe lattice, we can use self-similarity to compute the average cluster size. If $v$ is occupied, we can think of the cluster it belongs to as made up of the node $v$ itself plus the $B$ sub-clusters, which are at most $\theta$, its neighbors belong to. Combining this observation, the formula for the size of the perimeter of an $s$-cluster and Section A.1, we find the recursive relation
$\kappa_{\alpha}^{\text{Bethe}}=1+\theta B.$ (24)
Since each of the $\theta$ neighbors has probability $p$ of being occupied, and in that case, it has $\theta-1$ new neighbors that contribute to the cluster, we get
$B=p(1+(\theta-1)B)p.$ (25)
Via substitution, we find
$\kappa_{\alpha}^{\text{Bethe}}=\frac{1+p}{1-(\theta-1)p}.$ (26)
@@
As an example of this upper bound via the Bethe lattice, in Fig. 7, we plot cluster size distributions against the physical error rate $p$ in a circuit-level noise model for the $\llbracket 144,12,12\rrbracket$ code. The yellow distributions show the distribution of cluster sizes of the optimal factorization and the blue distributions are obtained from a BP-guided implementation of an LSD decoder (BP+LSD). The dashed line represents the average cluster size for the Bethe lattice as in Eq. 26 with parameter $\theta=139$.

### v2

@@ 3 $\mathbf{\lambda}$: fault node soft information vector
4 $\mathbb{I}\coloneqq[\,]$: list of invalid clusters $\mathrm{cl_{i}}$
5 $\mathbb{V}\coloneqq[\,]$: list of valid clusters $\mathrm{cl_{\ell}}$
@@
7 $\mathrm{cl_{i}}$ = create_cluster($s_{i}$)
8 $\mathbb{I}$.add($\mathrm{cl_{i}}$)
@@ 7 $\mathrm{cl_{i}}$ = create_cluster($s_{i}$)
14 for $\mathrm{cl_{i}},\mathrm{cl_{\ell}}\in\mathbb{I}\cup\mathbb{V}$ do
      // check if any clusters $\mathrm{cl_{i}},\mathrm{cl_{\ell}}$ collide
[…]
@@
[…]
18       $\mathbb{I}$.remove($\{\mathrm{cl_{i}},\mathrm{cl_{\ell}}\}$)
19       $\mathbb{V}$.remove($\{\mathrm{cl_{i}},\mathrm{cl_{\ell}}\}$)
20       $\mathbb{I}$.add($\mathrm{cl_{i\cup\ell}}$)
@@
24      valid = cl.check_validity($\mathbf{s}_{[\mathrm{cl}]}$)
@@
30 for $\text{cl}\in\mathbb{V}$  do
31   local_decodings.append(cl.plu_solve($\mathbf{s}_{[\mathrm{cl}]})$)
@@ 4.2 On-the-fly elimination
To grow and merge clusters, multiple smaller steps are necessary. As detailed above, these steps include identifying fault nodes/column indices of the decoding matrix $H$ by which the invalid clusters will grow and determining whether an added fault node will lead to two or more clusters merging into a single one – see Definition˜4.4. For simplicity, we first describe the case of sequential cluster growth. Our on-the-fly procedure can analogously be applied in a parallel implementation, see Appendix B.
Let $C_{i}$ be an active cluster, that is, $\left(H_{[C_{i}]},\mathbf{s}_{[C_{i}]}\right)$ does not define a solvable decoding problem as $s_{[C_{i}]}\notin\textsc{image}(H_{[C_{i}]})$. To grow cluster $C_{i}$, we consider candidate fault nodes $v_{j}\in\Lambda(C_{i})$ – fault nodes not already in $C_{i}$ but connected to check nodes on its boundary $\beta(C)$, see Eq.˜4. The candidate fault node with the highest probability of being in error according to the soft information vector $\mathbf{\lambda}\in\mathbb{R}^{n}$ is selected. Once $v_{j}$ has been chosen, we check whether its neighboring detector nodes are boundary nodes of any other (valid or invalid) clusters i.e., we check for collisions, cf Definition˜4.4. If this is not the case, we proceed as follows. We now assume that the active cluster $C_{i}$ described by sub-matrix $H_{[C_{i}]}$ that has a PLU factorization of the form
@@
where $P_{i},L_{i},U_{i}$ are as in Eq.˜6. Adding a fault node $v_{j}$ to the cluster is equivalent to adding a (sparse) column vector $\mathbf{b}$ to $H_{[C_{i}]}$, i.e.,
@@
$\begin{pmatrix}\begin{array}[]{c|c}U_{i}&L_{i}^{-1}P_{i}^{T}\mathbf{b}_{i}\\ 0&\mathbf{b}_{i*}\end{array}\end{pmatrix}$, (9)
where $\mathbf{b}_{i}$ is the projection of $\mathbf{b}$ onto the detectors/row coordinates that are enclosed by $C_{i}$. Similarly, $\mathbf{b}_{i*}$ is the projection onto detector coordinates not enclosed by $C_{i}$. Importantly, applying the operators $L_{i}^{-1}$ and $P_{i}^{T}$ does not affect the support of $\mathbf{b}_{i}$ and $\mathbf{b}_{i*}$, as both these operators act solely on the support of $C_{i}$. Combining this with Eq.˜9, we note that to complete the PLU factorization of $H_{[C_{i}\cup\{v_{j}\}]}$ only $\mathbf{b}_{i*}$ has to be reduced, which has a computational cost proportional to its weight – crucially only a small constant for bounded LDPC matrices $H$.
@@
Suppose that the selected fault node $v_{j}$ by which the cluster $C_{i}$ is grown is connected to a check node in the boundary $\beta(C_{\ell})$, with $C_{i}\neq C_{\ell}$. Let $\mathbf{b}$ be the column of $H$ associated with the fault node $v_{j}$. Re-ordering its coordinates if necessary, we can write $\mathbf{b}$ as $(\mathbf{b}_{i},\mathbf{b_{\ell}},\mathbf{b_{*}})$ where $\mathbf{b}_{i},\mathbf{b_{\ell}}$, and $\mathbf{b_{*}}$ are the projections of $\mathbf{b}$ on the row coordinates contained in $_{i}, CC_{\ell}$, and neither of them, respectively. Thus, using a block matrix notation, for the combined cluster $C_{i}\cup C_{\ell}\cup\{v_{j}\}$, we have
@@
$\begin{pmatrix}\begin{array}[]{c|c|c}U_{i}&0&L_{i}^{-1}P_{i}^{T}\mathbf{b}_{i}\\ \hline\cr 0&U_{\ell}&L_{\ell}^{-1}P_{\ell}^{T}\mathbf{b}_{\ell}\\ \hline\cr 0&0&\mathbf{b}_{*}\end{array}\end{pmatrix}$. (11)
@@
Here, by slight misuse of notation, we group the non-zero rows of $U_{m}$ in the index set $(m,\bullet)$, and its zero rows in the set $(m,\bot)$; we regroup the coordinates of vector $\mathbf{u}$ accordingly. We remark that the row sets $\bullet$ and $\bot$ are distinct from the row set * identified when writing $\mathbf{b}$ as the combination of its projection onto row coordinates enclosed by $C_{i}$ and outside it. By identifying the appropriate row sets for the clusters $_{i}, CC_{\ell}$ and the fault node $\{v_{j}\}$ as detailed in Eq.˜17, we can construct a block-swap matrix to bring $H_{[C_{i}\cup C_{\ell}\cup\{v_{j}\}]}$ into the form
$\begin{pmatrix}\begin{array}[]{c|c|c}U_{i,\bullet}&0&\mathbf{u}_{i,\bullet}\\ \hline\cr 0&U_{\ell,\bullet}&\mathbf{u}_{\ell,\bullet}\\ \hline\cr 0&0&\mathbf{u}_{i,\bot}\\ \hline\cr 0&0&\mathbf{u}_{\ell,\bot}\\ \hline\cr 0&0&\mathbf{b}_{*}\end{array}\end{pmatrix}$, (18)
and similarly for its PLU factors. Since the matrix in Eq.˜18 has the same form as the one in Eq.˜8, the algorithm can proceed from this point onward as in the case of the addition of a single fault node to a cluster. In conclusion, via a swap transformation, we can effectively reduce the problem of merging two clusters to the problem of adding a single fault node to one cluster.
@@ 4.3 Numerical decoding simulations
[…]
  • The measurement result of an XX/$Z$ basis measurement is flipped with probability $p$.
@@
The syndrome extraction circuits for the family of HGP codes presented in Appendix C.2 and results presented in Section 2.3 are obtained from the minimum edge coloration of the Tanner graphs associated to the respective parity check matrix, see Ref. [13] for details. In particular, we generate associated Stim files of $r=12$ noisy syndrome extractions using a publicly available implementation of the aforementioned coloration circuit by Pattison [57]. In this case, the standard circuit-level noise model described at the beginning of this section is employed. We decode $X$ and $Z$ detectors separately using a $(3,1)$-overlapping window decoder. That is, for each decoding round, the decoder obtains the detection events for $w=3$ syndrome extraction cycles and computes a correction for the entire window. However, it only applies the correction for a single $(c=1)$ syndrome extraction cycle, specifically the one that occurred the furthest in the past. For more details on circuit-level overlapping window decoding, see Ref. [58]. We have chosen $w=3$ as this was the value used in Ref. [13]. Note that it is possible that (small) decoding improvements could be observed by considering larger values of $(w,c)$ for the overlapping window decoder.
@@ 4.4 Parallel algorithm
We propose a parallel version of the LSD algorithm (P-LSD) in Appendix B that uses a parallel data structure, inspired by Refs. [60, 61], to minimize synchronization bottlenecks. We discuss the parallel algorithm in more detail in Appendix B. There, we derive a bound on the parallel depth of P-LSD, that is, roughly the maximum overhead per parallel resource of the algorithm. We show that the depth is $O(\operatorname{polylog}(n)+\kappa^{3})$ in the worst-case, where $n$ is the number of vertices of the decoding graph and $\kappa$ is the maximum cluster size. A crucial factor in the runtime overhead of P-LSD is given by the merge and factorization operations. We contain this overhead by (i) using the parallel union-find data structure of Ref. [60] for cluster tracking and (ii) using parallel on-the-fly elimination to factorize the associate matrices. If we assume sufficient parallel resources, the overall runtime of parallel LSD is dominated by the complexity of computing the decoding solution for the largest cluster. To estimate the expected overhead induced by the cluster sizes concretely, we (i) investigate analytical bounds and (ii) conduct numerical experiments to analyze the statistical distribution of clusters for several code families, see Appendix A.
@@ 7 Acknowledgments
This work was done in part while L.B. was visiting the Simons Institute for the Theory of Computing.
L.B. and R.W. acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 101001318) and Millenion, grant agreement No. 101114305). This work was part of the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus, and has been supported by the BMWK on the basis of a decision by the German Bundestag through project QuaST, as well as by the BMK, BMDW, and the State of Upper Austria in the frame of the COMET program (managed by the FFG).
J. R. is funded by an EPSRC Quantum Career Acceleration Fellowship (grant code: UKRI1224). J. R. further acknowledges support from EPSRC grants EP/T001062/1 and EP/X026167/1.
The Berlin team has been funded by BMBF (RealistiQ, QSolid), the DFG (CRC 183), the Munich Quantum Valley, the Einstein Research Unit on Quantum Devices, the Quantum Flagship (PasQuans2, Millenion), and the European Research Council (ERC DebuQC). For Millenion and the Munich Quantum Valley, this work is the result of joint-node collaboration.
@@ Appendix A Clusters on decoding graphs
To model an error distribution on the fault graph $\mathcal{F}(H)$, we assume that each fault node is occupied randomly with probability $p$ and empty with probability 1p1-p. Hence, for $n$ faults, we expect $np$ faults to be occupied, and n(1p)n(1-p) to be empty. This setup defines a site percolation problem, which involves studying the distribution of groups of neighboring occupied faults, called clusters. We call any instance of the site percolation problem an error defined on the fault graph $\mathcal{F}(H)$, and we expect clusters to have varying sizes and shapes depending on both the fault graph connectivity and the flip probability $p$.
@@
    the number $\nu$ of clusters of the error and
@@
More broadly, the expected performance depends on the average size of the clusters of the error.
In general, if we assume sufficient parallel resources, the LSD algorithm can use $\nu$ cores and run in $O(\kappa^{3})$ time in the worst case, where $\kappa$ represents the size of the largest cluster in a given instance. However, from a practical perspective, the largest cluster size for each sampled error (shot) is not the sole relevant figure of merit. As our statistical analysis suggests and our numerical investigation confirms, most errors exhibit minimal variance in their cluster size distributions, meaning that the clusters have the same size up to small fluctuations. Consequently, a useful practical proxy for the expected parallel runtime of LSD is the average cluster size per shot, $\kappa_{\alpha}$, rather than the worst-case cluster size.
Following this reasoning, we collect statistical data on these three quantities $\nu,\kappa$ and $\kappa_{\alpha}$, for a given noise model of interest. We study the expectation value of $\nu,\kappa$ and $\kappa_{\alpha}$ for:
@@
In Fig. 3, we report the statistics for the fault graph obtained from the circuit-level noise simulation of the $\llbracket 144,12,12\rrbracket$ code [12]. As we can see, for noise values below the threshold, there is no statistical difference between the two different cluster distributions, indicating that the clusters found by BP+LSD are of minimal size and hence (close to) optimal.
@@ A.1 Analytic cluster bounds
Figure 7: Distribution of the cluster size statistics of the $\llbracket 144,12,12\rrbracket$ bivariate bicycle code of Ref. [12] under circuit-level noise with strength pp. Markers show the mean of the distribution while shapes are violin plots of the distribution obtained from $10^{5}$ samples. Yellow distributions show statistics for the optimal factorization while the blue distributions show statistics for the factorization returned by BP+LSD. The dashed line represents the expected average cluster size for the Bethe lattice, $\kappa_{\alpha}$, for $\theta=139$. Markers and distributions are slightly offset from the actual error rate to increase readability.
@@
If $H$ is an LDPC matrix with constant row and column weight $r$ and $c$, respectively, the associated fault graph $\mathcal{F(H)}$ is a regular graph of degree $\theta=c(r-1)$, with ${\theta\cdot n}/{2}$ edges. Let $n_{s}^{p}(v)$ be the number of clusters of size $s$ containing the vertex $v$. In general, this takes the form
$p^{s}(1-p)^{|\mathrm{perimeter}|},$ (19)
where the perimeter of a cluster is the set of non-occupied nodes adjacent to at least one node in the cluster. The probability that the node $v$ belongs to a particular cluster of size $s$ is $n_{s}^{p}(v)s$, as $v$ can be any of the $s$ nodes in the clusters. On our percolation problem, by assumption, the probability that a given node is occupied is $p$. Moreover, if a node is occupied, it belongs to a cluster of size $\tilde{s}$ for some $\tilde{s}=1,\dots,n$, where $n$ is the number of nodes in the graph. Hence, for a given node in the graph, the probability $p$ of being occupied must equate the sum of the probabilities of belonging to a cluster of size $\tilde{s}=1,\dots,n$. Hence, for an arbitrary node in the graph, it holds that
$\sum_{s}n_{s}^{p}(v)s=p.$ (20)
That is, the probability $n_{s}^{p}(v)$ that a node belongs to an $s$-cluster times the number of nodes/choices for that cluster. A quantity of interest for decoding is an estimate of the size of the cluster a random node belongs to. In other words, the probability $w_{s}$ that a node is occupied and belongs to an $s$-cluster. Explicitly, using Equation˜20, we find
@@
and therefore the average cluster size is given by
@@
For the Bethe lattice, we can use self-similarity to compute the average cluster size $\kappa_{\alpha}^{\mathrm{Bethe}}$. If $v$ is occupied, we can think of the cluster it belongs to as made up of the node $v$ itself plus the sub-clusters i.e., subsets of connected vertices in that cluster, rooted in its occupied neighbors. Each occupied neighbor contributes to the cluster rooted in $v$ with itself (add 1) and the sub-clusters not rooted in $v$ itself, which at most $\theta-1$; by self-similarity, these sub-clusters have same average size as the cluster rooted in $v$, and contribute to the count adding at most $(\theta-1)\kappa_{\alpha}^{\mathrm{Bethe}}$ vertices. Thus, if the vertex $v$ is occupied, it belongs to a cluster of size $1+(\theta-1)\kappa_{\alpha}^{\mathrm{Bethe}}$. As each neighbor of the vertex $v$ is empty with probability (1p)(1-p), and occupied with probability $p$, we find
$\kappa_{\alpha}^{\mathrm{Bethe}}$ $=p\left(1+(\theta-1)\kappa_{\alpha}^{\mathrm{Bethe}}\right),$ (23)
and via substitution:
$\kappa_{\alpha}^{\text{Bethe}}=\frac{p}{1-p(\theta-1)}.$ (24)
@@
As an example of this upper bound via the Bethe lattice, in Fig.˜7, we plot cluster size distributions against the physical error rate $p$ in a circuit-level noise model for the $\llbracket 144,12,12\rrbracket$ code. The yellow distributions show the distribution of cluster sizes of the optimal factorization, and the blue distributions are obtained from a BP-guided implementation of an LSD decoder (BP+LSD), where the samples are post-selected on errors for which BP does not converge. That is, errors correctable by BP only are not included here for the BP+LSD distributions. The dashed line represents the average cluster size for the Bethe lattice as in Eq.˜24 with the parameter $\theta=139$.